Accelerator Update for October 1, 2008 - October 3, 2008
Wednesday Oct 1
The day shift began with Operations monitoring the Tevatron (TeV) store 6465 (185.7E30) and the Recycler stash (68.8E10), with the Antiproton Source (Pbar) stacking (22.3E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams.
At 8:31 AM, a MI RF station (MIRF5) tripped off. It reset.
At 8:54 AM, Operations reported that they were switching off the power for two power supplies (QP3 and HP3DS). An operator reported that the contacts for HP3DS did not open. The High Voltage Coordinator reported that the linkage for the knife switches had failed. He opened the disconnect in the upstream transformer so technicians could investigate further.
Operations prepared the F sector manholes for access and safety system testing at 1:24 PM.
A mechanical support expert reported at 3:47 PM that they had removed a considerable amount of the MTA - MuCool shield wall. The rest of the wall will be removed on Thursday.
Operations reported at 3:58 PM that the Recycler had stashed antiprotons five times during the day shift.
Operations terminated store 6465 at 4:18 PM.
Operations prepared D0 for an access at 4:34 PM. They came out 90 minutes later.
At 5:15 PM, the Pelletron tripped off. An expert investigated and got it restarted.
Operations put the TeV into shot setup at 7:07 PM.
Operations established store 6467 at 8:24 PM with an initial luminosity of 223.66E30.
Operations reported at midnight that the Recycler had stashed four times during the evening shift.
Thursday Oct 2
The midnight shift began with Operations monitoring store 6467 (129.62E30) and the stash (150.8E10), with Pbar stacking (21.2E10), and with the MiniBooNE and NuMI experiments taking beam.
At 3:06 AM, Operations reported that the Pbar Sweeper power supply tripped off for the first of two times in 18 minutes and the second trip took down the Pbar and NuMI beam permits.
Operations reported at 8 AM that the Recycler had stashed four times during the midnight shift.
Operations terminated store 6467 at 8:49 AM.
At 9:57 AM, a Linac RF station (LRF4) tripped off. It reset.
Cryo system technicians reported at 10:02 AM that they were replacing a power supply associated with the E1 wet engine.
At 10:49 AM, Operations reported Booster RF watchdog trips due to low LCW supply pressures. Experts and operators investigated.
Operations reported at 10:54 AM that they had put the TeV through a dry squeeze for the first of two times in about 30 minutes.
Operations put the Pbar into standby at 10:59 AM due to the Booster LCW problem.
Operations reported at 11:14 AM that the Booster LCW supply pressures were coming back up. A Booster expert reported that for some unknown reason the computer that controlled the pressure ordered the supply pumps to slow down. FESS personnel put all the pumps into manual operations. The investigation continued.
At 11:23 AM, LRF2 tripped off. It reset.
At 11:26 AM, a Linac Klystron RF station (KRF4) tripped off. It would not reset. An expert investigated.
Operations reported at 11:40 AM that the Booster RF system and kickers were back up and running.
Operations resumed sending beam to Pbar, NuMI, and MiniBooNE at 11:42 AM.
At 11:45 AM, BRF15 tripped off. It reset.
At 11:56 AM, BRF9 tripped off. It reset.
At 11:57 AM, Recycler experts reported that the Pelletron was down and would not reset. Experts investigated.
Operations put the TeV into shot setup at noon.
At noon, KRF4 tripped off for the first of three times in five minutes. It reset.
At 12:27 PM, BRF13 tripped off. It reset.
At 12:33 PM, MIRF5 tripped off. It reset.
At 1 PM, a Booster vacuum pump (IPRF17) tripped off. It reset.
At 1:21 PM, KRF2 tripped off for the first of four times in 20 minutes. It reset.
Operations established store 6469 at 1:29 PM with an initial luminosity of 185.03E30.
At 2:05 PM, the NuMI beam permit tripped off when the current of a power supply (E:QD108) went low. Everything reset.
At 2:08 PM, MIRF5 tripped off. It reset.
Operations prepared MI-31 for access by Pelletron experts at 2:31 PM.
Operations reported at 3:09 PM that the Pbar stack was at 150E10. Since the Electron Cooling system was down and cannot take a stash, a Pbar expert began to tune. (Later, the Pelletron experts reported that the Pelletron’s Generator Volt Meter had a mechanical failure. The Recycler will resume stashing without the Pelletron. The repairs should be complete by the end of the shutdown next week.)
Operations reported at 3:35 PM that they were going to a NuMI only run mode.
At 5:27 PM, MIRF5 tripped off. It reset.
Operations reported at 5:48 PM that they had completed a stash to the Recycler and that Pbar had resumed stacking.
At 8:13 PM, a Booster vacuum pump (IPRF23) was tripping off and causing BRF14 to trip off. Everything reset.
Operations reported at 9:59 PM that the Pbar stack reached 150E10. Operations put Pbar into standby and went into a NuMI only run mode.
Operations reported at midnight that the Recycler had stashed once during the evening shift. (Due to the problem with the Pelletron there will be fewer, but larger stashes from Pbar.)
Friday Oct 3
The midnight shift began with Operations monitoring store 6469 (60.79E30) and the stash (121.2E10), with Pbar stacking (149.2E10), and with the MiniBooNE and NuMI experiments taking beam.
At 6:36 AM, KRF6 tripped off. It reset.
TeV experts began a study at 7:03 AM.
The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store. A weeklong shutdown of the accelerators will begin on Sunday, 10/5/08.
Comments and Suggestions
|last modified 10/03/2008 email Fermilab|
|Security, Privacy, Legal|